Calculate the approximate force on a square meter of sail, given the horizontal velocity of the wind is 9.16 m/s parallel to its front surface and 4.9 m/s along its back surface. Take the density of air to be 1.29 kg/m3.

Round your answer to 2 decimal places


1
Expert's answer
2021-05-03T10:29:45-0400

"\text{force per square meter is the definition of pressure}".

"\text{the pressure on the sail will be the difference in pressure}\newline
\text{ on the inside and outside of the sail}"

"\text{according to bernoulli law:}"

"\frac{\rho v^2}{2}+\rho gh+P= const"

"\text {for front surface:}"

"\frac{\rho*9.16^2}{2}+\rho gh+P_1"

"\text {for back surface:}"

"\frac{\rho*4.9^2}{2}+\rho gh+P_2"

"\frac{\rho*9.16^2}{2}+\rho gh+P_1 =\frac{\rho*4.9^2}{2}+\rho gh+P_2"

"P_2-P_1=\frac{\rho*9.16^2}{2}-\frac{\rho*4.9^2}{2}=\frac{1.29*(9.16^2-4.9^2)}{2}\approx29.95"

"\text{Answer: }29.95"